OpenStudy (anonymous):

find the missing value to the nearest hundreth sin______ =9/20

OpenStudy (anonymous):

can someone please help me?

OpenStudy (anonymous):

@saynabahmed13 can u help me?

OpenStudy (anonymous):

or @lizz123

OpenStudy (anonymous):

anyone?

OpenStudy (lizz123):

Here's what I would do without a calculator or a knowledge of the arcsin function. It's a crude kind of what's called a numerical method. Suppose I wanted to solve sin x = 7/16. 7/16 = .4375 I'd make a guess: sin 45(degrees) = .707 (make a smaller guess) sin 20 = .342 (make bigger guess) sin 30 = .5 (go between 20 and 30) sin 25 = .423 (go between 25 and 30) sin 26 = .438 ... and so on.

OpenStudy (lizz123):

webite found @: http://openstudy.com/study#/updates/5356bf76e4b01d52d929ecec

OpenStudy (lizz123):

@hi_bye

OpenStudy (anonymous):

thank you!

OpenStudy (lizz123):

yw

OpenStudy (anonymous):

how do u give someone a medal on here?

OpenStudy (lizz123):

best response

OpenStudy (lizz123):

click on that

OpenStudy (anonymous):

well there u go! i gave u a medal for your help! thank you very much!

OpenStudy (lizz123):

no prob gave u 1 2

OpenStudy (anonymous):

so to figure out the sin of 9/20 i am going to do 9/20=.45 but then what so i do?

OpenStudy (anonymous):

i mean 9/16=.5625 so what do i do now?

OpenStudy (anonymous):

@lizz123